# GPU-Accelerated Vecchia Approximations of Gaussian Processes for Geospatial Data using Batched Matrix Computations

### v1

@@ IV-A Gaussian Process and Likelihood Function
Statistical modeling involves analyzing correlations among various spatial or spatio-temporal points distributed regularly or irregularly over a geographical area. In the purely spatial scenario, these datasets are assumed to be realizations observed from a Gaussian spatial random field. Assume there are $n$ locations labeled from 1 through $n$, as in ${\bf s}_{1},\ldots,{\bf s}_{n}\in\mathbb{R}^{d}$. Let $y_{i}:=y\left({\bf s}_{i}\right)\in\mathbb{R}$ represents the observation associated with location ${\bf s}_{i}$ and denote the vector $\mathbf{y}=\left(y_{1},\ldots,y_{n}\right)^{\top}$. In a Gaussian process, the data $\mathbf{y}$ are modeled as $\mathbf{y}\sim\mathcal{N}\left(\bm{\mu},\bm{\Sigma}_{\bm{\theta}}\right)$, where $\bm{\Sigma}_{\bm{\theta}}$ is a covariance matrix with $(i,j)$ entry determined by a given covariance function $C_{\bm{\theta}}\left({\bf s}_{i},{\bf s}_{j}\right)$ relying on a vector of covariance parameters, ${\bm{\theta}}$. Without loss of generality, we assume that $\mathbf{y}$ has a mean of zero. Statistical inference about $\bm{\theta}$ is often based on the Gaussian log-likelihood function, which entails a cubic computational complexity [5]:
@@ V-B Batched Vecchia Approximated Likelihood Algorithm
Having a set of geospatial data consisting of $n$ spatial locations and their corresponding observations denoted by $\mathbf{y}$ and considering any permutation (reordering) $\tau$ of the given locations, the likelihood, or the joint density, for the observations $\mathbf{y}$ can be written as a product of conditional densities as follows:
@@
$=p_{\bm{\theta}}\left(y_{1}^{\tau}\right)\prod_{i=2}^{n}p_{\bm{\theta}}\left(y_{i}^{\tau}\mid y_{1}^{\tau},\ldots,y_{i-1}^{\tau}\right),$ (7)
reflecting the invariance of the joint density to arbitrary relabeling of the observations. Vecchia approximation replaces the complete conditioning vectors $\left(y_{1}^{\tau},\ldots,y_{i-1}^{\tau}\right)$, with a subvector. Specifically, let $\left\{j_{i1},\ldots,j_{im_{i}}\right\}$ be a set of integers between 1 and $i-1$, $\mathbf{y}^{\tau}_{J_{i}}=(y_{j_{i1}}^{\tau},\ldots,y_{j_{im_{i}}}^{\tau})^{\top}$ and define the approximation
@@
Here, we define $J$ as a collection of sets, represented as $J=\left\{J_{1},\ldots,J_{n}\right\}$, where each $J_{i}$ is defined as $\left\{j_{i1},\ldots,j_{im_{i}}\right\}$ and is referred to as the conditioning set containing $m_{i}$ nearest neighbors for observation $i$. In the Vecchia algorithm, the approximation quality relies on selecting $J$ since the approximate likelihood deviates from the exact likelihood when $m_{i}\neq i-1$ for any observation $i$. Additionally, the approximation quality is influenced by the permutation $\tau$ as discussed in [21].

### v3

@@ IV-A Gaussian Process and Likelihood Function
Statistical modeling involves analyzing correlations among various spatial or spatio-temporal points distributed regularly or irregularly over a geographical area. In the purely spatial scenario, these datasets are assumed to be realizations observed from a Gaussian spatial random field. Consider a set of $n$ locations, ${\bf s}_{1},\ldots,{\bf s}_{n}\in\mathbb{R}^{d}$, and their observations, $\mathbf{y}=\left(y_{1},\ldots,y_{n}\right)^{\top}$ which is indicated by $y_{i}:=y\left({\bf s}_{i}\right)\in\mathbb{R}$. We model the data $\mathbf{y}$ with Gaussian process, $\mathbf{y}\sim\mathcal{N}\left(\bm{\mu},\bm{\Sigma}_{\bm{\theta}}\right)$, where $\bm{\Sigma}_{\bm{\theta}}$ is a covariance matrix with $(i,j)$ entry determined by a given covariance function $C_{\bm{\theta}}\left({\bf s}_{i},{\bf s}_{j}\right)$ relying on a vector of covariance parameters, ${\bm{\theta}}$. Without loss of generality, we assume that $\mathbf{y}$ has a mean of zero. Statistical inference about $\bm{\theta}$ is often based on the Gaussian log-likelihood function, which entails a cubic computational complexity [5]:
@@ V-B Batched Vecchia Approximated Likelihood Algorithm
In an analysis involving a set of geospatial data points $\mathbf{y}$ across various locations, and their observed values, when we apply any reordering $\tau$ to the sequence of these locations, the likelihood expression for the dataset retains its form when represented as a sequence of conditional densities,
@@
$=p_{\bm{\theta}}\left(y_{1}^{\tau}\right)\prod_{i=2}^{n}p_{\bm{\theta}}\left(y_{i}^{\tau}\mid y_{1}^{\tau},\ldots,y_{i-1}^{\tau}\right).$ (7)
Vecchia approximation is expressed as (9) and it replaces the conditioning vectors $\left(y_{1}^{\tau},\ldots,y_{i-1}^{\tau}\right)$ with $(y_{j_{i1}}^{\tau},\ldots,y_{j_{im_{i}}}^{\tau})$ which are nearest neighbors to the target $y_{j_{i}}^{\tau}$; $J_{i}$ is defined as $\left\{j_{i1},\ldots,j_{im_{i}}\right\}$ and $J=\left\{J_{1},\ldots,J_{n}\right\}$. [21]
@@
The accuracy of the Vecchia algorithm’s approximation is dependent on the choice of permutation $\tau$ and $J$, because the permutation $\tau$ creates specific ordering from 2D to 1D and the $J$ represents the deviations from the exact likelihood. [21] [13]